OpenStudy (anonymous):

ok so it asks.... "The size of an exponentially growing bacteria colony doubles in 8 hours. How long will it take for the number of bacteria to triple? Give your answer in exact form and decimal form."

OpenStudy (campbell_st):

so your model is \[P = P_{0}e^{kt} \]

OpenStudy (anonymous):

yes i believe it is. but i'm missing so many variables how do i solve it without\[P _{o}\] and P @campbell_st

OpenStudy (campbell_st):

ok... so the initial population is 1 so it doubles to 2 then t = 8 you need to find the growth constant so \[2 = 1 \times e^{8k}\] solve for k by taking the ln of both sides \[\ln(2) = 8k\] when you know the value of k set P to 3 and solve for t... does that make sense?

OpenStudy (anonymous):

Oh snap! nice!

OpenStudy (anonymous):

but wait when i have \[3=e ^{\ln(2)/8\times(t)}\] how do i bring it down and solve. that t confuses me

OpenStudy (campbell_st):

thats the way I see the question.. so take the log of both sides... \[\ln(3) = \frac{\ln(2)}{8} \times t\]

OpenStudy (campbell_st):

thats why they want an exact answer and a decimal answer...

OpenStudy (anonymous):

holy cow! brilliant! thanks!